### v3

# GrEmLIn: A Repository of Green Baseline Embeddings for 87 Low-Resource Languages Injected with Multilingual Graph Knowledge

Abstract
Contextualized embeddings based on large language models (LLMs) are available for various languages, but their coverage is often limited for lower resourced languages. Using LLMs for such languages is often difficult due to a high computational cost; not only during training, but also during inference. Static word embeddings are much more resource-efficient ("green"), and thus still provide value, particularly for very low-resource languages. There is, however, a notable lack of comprehensive repositories with such embeddings for diverse languages. To address this gap, we present GrEmLIn, a centralized repository of green, static baseline embeddings for 87 mid- and low-resource languages. We compute GrEmLIn embeddings with a novel method that enhances GloVe embeddings by integrating multilingual graph knowledge, which makes our static embeddings competitive with LLM representations, while being parameter-free at inference time. Our experiments demonstrate that GrEmLIn embeddings outperform state-of-the-art contextualized embeddings from E5 on the task of lexical similarity. They remain competitive in extrinsic evaluation tasks like sentiment analysis and natural language inference, with average performance gaps of just 5-10% or less compared to state-of-the-art models, given a sufficient vocabulary overlap with the target task, and underperform only on topic classification. Our code and embeddings are publicly available at https://huggingface.co/DFKI [^1] All vectors are available on Huggingface as single model pages. Each page starts with DFKI/glove..
GrEmLIn: A Repository of Green Baseline Embeddings for 87 Low-Resource Languages Injected with Multilingual Graph Knowledge
Daniil Gurgurov1  Rishu Kumar1  Simon Ostermann1,2 1German Research Center for Artificial Intelligence (DFKI) 2Centre for European Research in Trusted AI (CERTAIN) {daniil.gurgurov, rishu.kumar, simon.ostermann}@dfki.de
1 Introduction
Word embedding methods have revolutionized natural language processing (NLP) by capturing semantic relationships between words using co-occurrence statistics from large text corpora Mikolov et al. (2013a); Pennington et al. (2014); Bojanowski et al. (2017). This data-driven approach has significantly improved performance across numerous NLP tasks Lample et al. (2017); Xie et al. (2018); Almeida and Xexéo (2019).
While contextual representations like the ones based on BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and GPT Radford et al. (2019) nowadays provide better performance than static embeddings in many tasks, their training is computationally expensive Strubell et al. (2019); Bommasani et al. (2021) and ineffective for data-scarce languages due to their data hunger and the curse of multilinguality Conneau et al. (2020). Some approaches of efficient adaptation of especially large language models (LLMs) to languages other than English have been investigated in recent years Pfeiffer et al. (2020); Vykopal et al. (2024). However, even such approaches still require hardware during runtime, as embeddings need to be computed based on a forward pass for each new text that is processed. This is often prohibitive in low-resource (hardware) scenarios, and inefficient in terms of energy use. Also, such approaches are often not tailored to low-resource languages.
In contrast, static word embeddings continue to play a crucial role in specific tasks such as bias detection and removal Gonen and Goldberg (2019); Manzini et al. (2019), explaining word vector spaces Vulić et al. (2020b); Bommasani et al. (2020), and information retrieval Yan et al. (2018). Static word embeddings have the advantage of being parameter-free at inference time, as no neural network needs to be loaded for computing such representations; just a dictionary lookup is required. This makes them both attractive for low-resource hardware scenarios, and much more environment-friendly Strubell et al. (2019); Dufter et al. (2021). Existing resources for multingual embedding data bases Ferreira et al. (2016); Bojanowski et al. (2017); Grave et al. (2018a) often suffer from limited scope and outdated data, potentially worsening their ability to capture the dynamic nature of language and adequately support low-resource languages. We want to fill this gap by providing GrEmLIn, a large database of static word embeddings for 87 mid- and low-resource languages.
As for LLMs, the training of word embeddings suffers from the lack of high-quality data in low-resource languages (to a smaller degree). Incorporating other types of data for improving word representations is thus beneficial especially for low-resource languages. Knowledge graphs provide such an alternative to textual knowledge, with rich semantic and multilingual sources of information, including synonyms, antonyms, morphological forms, definitions, etimological relations, translations, and more Miller (1995); Navigli and Ponzetto (2012); Speer et al. (2017). Such structured and cross-lingual information can be used to improve the quality of classical word representations Faruqui et al. (2014); Sakketou and Ampazis (2020), which are only trained on co-occurence statistics.
To that end, we propose a new simple yet effective method for including graph information into word embeddings based on Mikolov et al. (2013b). We learn a projection matrix to map static embeddings to a combined space, effectively overcoming the limitations of retrofitting approaches that only enhance a limited vocabulary. This method combines the strengths of traditional word embeddings with the structured, multilingual information from knowledge graphs, resulting in more accurate and informative representations.
In summary, our contributions in this work are two-fold: First, we present GrEmLIn, a centralized resource of static word embeddings for 87 mid- and low-resource languages, specifically focusing on word embeddings trained with GloVe Pennington et al. (2014). Second, we propose an effective method to improve embeddings by incorporating more knowledge in the form of multilingual knowledge graphs, which is especially important for low-resource languages, where resources are usually very scarce. Our code is publicly available on GitHub [^2] https://github.com/d-gurgurov/GrEmLIn-Green-Embeddings-LRLs.
2 Related Work
We briefly describe the most prominent graph knowledge sources, word embeddings, and existing methods for improving embeddings with graphs.
Graph knowledge sources.
Among most used knowledge graphs for natural language are WordNet Miller (1995) and BabelNet Navigli and Ponzetto (2012). WordNet is a lexical database that organizes English words into sets of synonyms called synsets, providing short definitions and usage examples. BabelNet is a multilingual encyclopedic dictionary and semantic network, which integrates lexicographic and encyclopedic knowledge from WordNet, Wikipedia, etc., focused on named entities. In our work, we use ConceptNet Speer et al. (2017), a multilingual, domain-general knowledge graph that connects words and phrases from various natural languages with labeled, weighted edges representing relationships between terms. Unlike other knowledge graphs, ConceptNet is not a monolingual collection of named entities but focuses on commonly used words and phrases across multiple languages.
Word embeddings.
Word2Vec Mikolov et al. (2013a) uses shallow neural networks to produce word vectors. It comes in two types: Continuous Bag of Words (CBOW) and Skip-gram. CBOW predicts a word given its context, while Skip-gram predicts the context given a word. GloVe (Global Vectors for Word Representation) Pennington et al. (2014) word embeddings are created by aggregating global word-word co-occurrence statistics from a corpus. The resulting vectors capture both local and global semantic relationships. FastText Bojanowski et al. (2017) extends Word2Vec by representing words as bags of character n-grams, capturing subword information and handling out-of-vocabulary words more effectively. FastText is particularly useful for morphologically rich languages. Numberbatch, part of the ConceptNet project Speer et al. (2017), is a set of word embeddings that integrates knowledge from ConceptNet with distributional semantics from GloVe and Word2Vec. Numberbatch uses a retrofitting approach Faruqui et al. (2014) to enhance embeddings with structured semantic knowledge. Retrofitting often results in a limited vocabulary for underrepresented languages Speer and Lowry-Duda (2017) since the retrofitting process relies on existing semantic relationships in ConceptNet to adjust the original embeddings.
Improving Embeddings with Knowledge Graphs.
There are various methods to improve word embeddings by incorporating external knowledge graphs or semantic networks Dieudonat et al. (2020). Retrofitting Faruqui et al. (2014) is a post-processing technique that adjusts pre-trained word embeddings using information from knowledge graphs or semantic lexicons. The key idea is to infer new vectors that are close to their original embeddings while also being close to their neighbors in the graph or lexicon. This is achieved by minimizing an objective function that balances the distance between the new vectors and the original embeddings, as well as the distance between connected nodes. Expanded retrofitting Speer et al. (2017), used for ConceptNet Numberbatch, optimizes over a larger vocabulary including terms from the knowledge graph not present in the original embeddings, but it still does not retrofit all the words in the original embedding space. Other existing methods that integrate contextualized embeddings with knowledge graph embeddings often use attention mechanisms, as demonstrated by works such as Peters et al. (2019), Zhang et al. (2019), and Gurgurov et al. (2024). These methods specifically enhance BERT embeddings by incorporating external knowledge bases.
3 Method
We propose a method for merging GloVe embeddings with graph-based embeddings derived from ConceptNet knowledge, while preserving the vocabulary size of GloVe, following two steps: First, we use singular value decomposition (SVD) Eckart and Young (1936) on concatenated word embeddings from GloVe and pointwise mutual information (PMI) based graph embeddings Speer et al. (2017) to generate a shared embedding space. We do so for the part of the vocabulary that is shared between GloVe and the knowledge graph. Second, we learn a linear transformation from GloVe into this joined space to obtain embeddings for all words in the original GloVe vocabulary.
3.1 GloVe Embeddings
We train GloVe embeddings using the original code. The model is trained by stochastically sampling nonzero elements from the co-occurrence matrix over 100 iterations, to produce 300-dimensional vectors. We use a context window of 10 words to the left and 10 words to the right. Words with fewer than 5 co-occurrences are excluded for languages with over 1 million tokens in the training data, and the threshold is set to 2 for languages with smaller datasets. We use data from CC100333https://huggingface.co/datasets/cc100 Wenzek et al. (2020); Conneau et al. (2020) for training the static word embeddings. We set $x_{max}=100$, $\alpha=\frac{3}{4}$, and use AdaGrad optimization Duchi et al. (2011) with an initial learning rate of 0.05.
3.2 Graph Embeddings
To build ConceptNet-based word embeddings, we follow the method used for constructing ConceptNet Numberbatch embeddings Speer et al. (2017). We represent the ConceptNet graph as a sparse, symmetric term-term matrix, where each cell is the sum of the occurences of all edges connecting the two terms. Unlike the original method, we do not discard terms connected to fewer than three edges, as we deal with low-resource languages.
We calculate embeddings from this matrix by applying pointwise mutual information (PMI) with context distributional smoothing of 0.75, clipping negative values to yield positive PMI (PPMI), which follows practical recommendations by Levy et al. (2015). We then reduce the dimensionality to 300 using truncated SVD and combine terms and contexts symmetrically to form a single matrix of word embeddings, called ConceptNet-PPMI. This matrix captures the overall graph structure of ConceptNet.
We compute ConceptNet-PPMI embeddings for the entire ConceptNet, covering 304 languages, which we call PPMI (All). Further, we construct separate graph embedding spaces, PPMI (Single), for each language, using only the portion of ConceptNet for that language. This approach is adopted since the initial co-occurence matrices for individual languages are less sparse while still being multilingual in nature.
3.3 Singular Value Decomposition (SVD)
We first concatenate GloVe and PPMI vectors for all words that are in the shared vocabulary, resulting in 600-dimensional vectors [^4] PPMI embeddings are normalized to be in the range of the Glove embeddings. Afterwards, we reduce the dimensionality and remove some of the variance coming from redundant features. The matrix $M$ representing merged GloVe and ConceptNet-PPMI can be approximated with a truncated SVD:
$M\approx U\Sigma V^{T}$
where $\Sigma$ is truncated to a $k^{\prime}\times k^{\prime}$ diagonal matrix of the $k^{\prime}$ largest singular values, and $U$ and $V$ are correspondingly truncated to have only these $k^{\prime}$ columns. $U$ is then used as a matrix mapping the original vocabulary to a smaller set of features [^5] We dismiss the weighting of $U$ by the singular values from $\Sigma$, which was noted to work better for semantic tasks Levy et al. (2015).
3.4 Linear Transformation
To obtain embeddings for the entire vocabulary from the original GloVe embedding space (i.e. not only the common words), we find a linear projection matrix between the spaces and project the GloVe embeddings onto the merged embedding space, similar to Mikolov et al. (2013c), using a gradient descent optimization on a linear regression model.
[…]
$W$ can be learned by solving the following optimization problem:
$\min_{W}\sum_{i=1}^{n}\|Wx_{i}-z_{i}\|^{2}$
which we solve as a linear regression problem with stochastic gradient descent optimization.
The resulting projection matrix is used to project the GloVe embeddings onto the merged embedding space.
4 Experiments
In this section, we describe the selected languages, tasks, and experiments conducted to evaluate the effectiveness of our proposed method.
4.1 Languages
We trained GloVe embeddings for 87 languages from the CC100 dataset Wenzek et al. (2020), focusing on languages categorized as low-resource (class 0 to 3) based on Joshi et al.’s classification (2020). For 72 of these languages, present in both CC100 and ConceptNet, we generated additional graph embeddings. The merging process involves enhancing the original GloVe embeddings with graph knowledge via SVD-reduced PPMI integration. Further details about these languages, including common vocabulary size between GloVe and ConceptNet, can be found in Part C of the Appendix.
4.2 Evaluation Data
We assess the embeddings using both intrinsic and extrinsic evaluation tasks. The intrinsic evaluation is performed using the MultiSimLex dataset Vulić et al. (2020a), which provides manually annotated data on semantic similarity consisting of 1888 examples across 12 languages, 4 of which overlap with our work. This task focuses on measuring the strength of similarity between word pairs (e.g., "lion – cat") independently of relatedness, making it a good test for how well embeddings capture semantic similarity.
For extrinsic evaluations, we focus on three downstream NLP tasks: Sentiment Analysis (SA), Topic Classification (TC), and Natural Language Inference (NLI). Due to the limited availability of intrinsic datasets for most low-resource languages, we prioritize these tasks to reflect real-world use cases, where high-quality word embeddings are crucial.
For SA, we compile data for 23 languages from multiple open sources, prioritizing mid- and low-resource languages for broader coverage across typological families. The details of these data sources are listed in Table 7 in the Appendix. Some datasets, such as those for Swahili, Nepali, Uyghur, Latvian, Slovak, Slovenian, Uzbek, Bulgarian, Yoruba, Bengali, Hebrew, and Telugu, are highly imbalanced in terms of class distribution. To mitigate this, we apply random undersampling to create a balanced version of the datasets. This step allows for a more robust comparison of the embeddings’ performance in low-data settings.
We further evaluate the embeddings on the TC task using the SIB-200 dataset Adelani et al. (2024). It offers multilingual data for topic classification, covering 200 languages, and was specifically designed to improve natural language understanding for under-resourced languages. Our experiments cover 57 languages, chosen based on their availability in both ConceptNet and CC100. The task is framed as a multi-label classification with the data distributed along 7 different classes. The dataset provides predefined train, validation and test splits, which consist of 701, 99, and 204 examples, respectively.
Lastly, we evaluate the embeddings on the NLI task, using the XNLI dataset Conneau et al. (2018). The XNLI dataset provides multilingual NLI examples for 15 languages, and for our experiments, we selected 5 of these languages: Swahili, Urdu, Greek, Thai, and Bulgarian. These languages were selected based on the availability of our GloVe, PPMI-enhanced embeddings, and the NLI dataset. Evaluating embeddings on the NLI task tests their ability to understand logical relationships between sentence pairs, an important capability for higher-level NLU tasks. Due to the simplicity of our models, we only utilize validation and test splits, consisting of 2,490 and 5,010 examples, respectively, for training and testing, excluding the original training split of nearly 400,000 examples.
4.3 Experimental Setup
We evaluate the embeddings using a Support Vector Machine (SVM) classifier Boser et al. (1992) for all extrinsic tasks—SA, TC, and NLI—reporting macro-averaged F1 scores for fair comparison. For the intrinsic MultiSimLex task, we use Spearman’s Rank Correlation Spearman (1961) to assess how well the embeddings’ similarity predictions align with human annotations.
For extrinsic tasks, sentence representations are constructed by summing word embeddings, which is a standard approach in NLP Mikolov et al. (2013d); Bowman et al. (2015); Williams et al. (2018), and then used as input features for the SVM. The SVM model is trained with a Radial Basis Function (RBF) kernel, which is commonly used for nonlinear classification problems. The regularization parameter $C$ is fixed at 1 for GloVe-based embeddings, balancing the trade-off between maximizing the margin and minimizing classification errors. This setup minimizes the impact of hyperparameters on the resulting scores.
For the NLI task, sentence representation follows the same method as above, but with an added step. We concatenate the sentence embeddings of the two input sentences (premise and hypothesis) to form the final input representation for the SVM. This approach enables the model to capture the relationship between the two sentences.
As for baselines, we use three strong pre-trained models:
  • FastText Grave et al. (2018b), a word embedding model that extends the traditional skip-gram model by representing words as bags of character n-grams, allowing it to effectively handle out-of-vocabulary words.
  • XLM-R-base Conneau et al. (2020), a transformer-based multilingual model. We obtain sentence embeddings by summing the model’s last hidden states.
  • E-5-base Wang et al. (2024), a state-of-the-art multilingual sentence embedding model known for its strong performance in multilingual tasks. This serves as a high-quality benchmark for our comparisons.
Cov.     ISO Contextualized Static
    E-5-B X-B     FT G GP(S) GP(A)
>90%     et .19 .03     .447 .341 .452 .422
    he .218 .057     .426 .336 .436 .429
[…]
<90%     cy .112 .039     .346 .276 .366 .357
    sw .212 .011     .408 .24 .319 .324
[…]
    All avg. .183 .034     .407 .298 .393 .383
Table 1: Spearman’s correlation scores on MultiSimLex across 4 languages, for E-5-B, XLM-R-B, FastText, GloVe (G), GloVe + PPMI (GP), Single and All, sorted by GloVe vocabulary coverage. The horizontal solid line separates languages with over 90% coverage (above) from those with less (below). Bold numbers indicate the maximum per line, for static and contextualized.
For the XLM-R-base and E-5 embeddings, we adjust the regularization parameter $C$ to 100. This adjustment accounts for the higher dimensionality of these embeddings, as lower $C$ values constrain their performance.
5 Results
We distinguish between static and contextualized embeddings by first comparing the static embeddings against each other, and then comparing them to the contextualized ones. The results from E-5-B are provided for reference but cannot be directly compared to our static embeddings due to the reasons outlined in Section 6.
5.1 Semantic Similarity
We evaluate the performance of the embeddings on the lexical semantics task using the MultiSimLex dataset, focusing on 4 languages: Estonian, Welsh, Swahili, and Hebrew.
As shown in Table 4.3, the GloVe+PPMI (Single) embeddings achieve the highest correlation scores for 3 out of 4 languages, demonstrating their ability to capture semantic similarities. For Swahili, FastText achieves the best result, although GloVe+PPMI remains competitive. In contrast, contextual embeddings such as XLM-R-base struggle in this intrinsic evaluation task, achieving lower correlation scores across all languages, which supports Vulić et al. (2020a). E-5 performs better than XLM-R but does not surpass the best-performing static embeddings.
These results underscore the continued relevance of graph-enhanced static embeddings in lexical semantic tasks, particularly for low-resource languages where training data may be scarce.
5.2 Sentiment Analysis
We evaluate the performance of the proposed GloVe+PPMI embeddings on the SA task for 23 mid- and low-resource languages. Table 5.2 presents the results for this task. Our findings show that both GloVe+PPMI (Single) and GloVe+PPMI (All) embeddings consistently outperform the original GloVe embeddings across most languages. GloVe+PPMI (Single) improves performance for 19 out of 23 languages, while GloVe+PPMI (All) improves results for 18 out of 23 languages when compared to GloVe.
When comparing GloVe with FastText embeddings, we observe that GloVe outperforms FastText in 12 out of 23 languages, with some languages showing comparable results.
Cov.     ISO Contextualized Static
    E-5-B X-B     FT G GP(S) GP(A)
>90%     ka .9 .845     .855 .861 .87 .861
    sl .881 .832     .743 .749 .779 .788
    ro .926 .872     .803 .805 .85 .847
[…]
>80%     he .929 .811     .782 .788 .824 .822
    si .895 .831     .846 .848 .85 .857
    sw .773 .665     .697 .68 .701 .714
    ug .881 .61     .792 .746 .811 .811
    lv .801 .74     .749 .783 .787 .787
    te .854 .831     .798 .806 .808 .817
    sk .911 .854     .73 .756 .806 .805
    mr .912 .886     .888 .903 .905 .902
    bg .884 .721     .793 .786 .801 .805
    mk .817 .736     .682 .716 .711 .7
[…]
<80%     su .855 .829     .805 .798 .822 .812
    am .861 .782     .815 .881 .86 .88
    ne .666 .519     .666 .643 .674 .688
    da .972 .927     .895 .863 .908 .903
    uz .858 .807     .822 .808 .806 .806
    bn .938 .837     .889 .875 .881 .878
    ur .818 .757     .678 .676 .746 .745
    az .787 .762     .75 .744 .746 .745
    cy .834 .795     .798 .77 .789 .801
    yo .764 .634     .696 .721 .709 .738
[…]
    All avg. .857 .778     .781 .783 .802 .805
Table 2: Macro Average F1 Scores for Sentiment Analysis per language, sorted by GloVe vocabulary coverage. Horizontal solid lines indicate 90% and 80% coverage by GloVe. Bold numbers indicate the maximum per line, for static and contextualized.
In contrast, XLM-R-base performs better than all static embedding configurations for 9 out of 23 languages, and E-5 outperforms most static embedding variants. While this underscores the power of contextualized models, the enhanced GloVe+PPMI embeddings remain competitive, with a drop of only 5% in performance, especially in low-resource settings. This suggests that static embeddings, when enriched with multilingual graph knowledge, remain competitive and provide a lightweight and efficient zero-parameter alternative for resource-constrained environments.
Cov.     ISO Contextualized Static
    E-5-B X-B     FT G GP(S) GP(A)
>80%     bg .563 .465     .465 .441 .481 .477
    el .546 .455     .484 .456 .496 .488
    sw .539 .437     .471 .438 .466 .468
[…]
<80%     ur .540 .472     .412 .44 .473 .471
    th .538 .461     .275 .284 .292 .3
[…]
    All avg. .545 .458     .421 .412 .442 .441
Table 3: Macro Average F1 Scores for Natural Language Inference per language, sorted by GloVe vocabulary coverage. The horizontal solid line separates languages with over 80% coverage (above) from those with less (below). Bold numbers indicate the maximum per line, for static and contextualized.
5.3 Natural Language Inference
In the NLI task using the XNLI dataset, we again observe consistent improvements in performance with the enhanced GloVe embeddings (Table 5.2). While GloVe outperforms FastText for only 2 out of 5 languages, the use of PPMI (Single) and PPMI (All) results in better performance for all 5 languages.
In comparison, XLM-R performs better than the static embedding variants for 1 out of the 5 languages, and E-5 outperforms all models in all languages. While transformer models like XLM-R excel in capturing complex semantic relationships between sentences, the performance of GloVe+PPMI remains competitive, with a drop of only 6% given a sufficient vocabulary overlap, especially in improving sentence-level reasoning and inference capabilities in low-resource languages.
5.4 Topic Classification
The results of the topic classfication task using the SIB-200 dataset are the only results where the contextualized models seem to have a clear advantage. The drop for a vocabulary coverage of over 95% is only at 10%, but over all languages the drop averages at 20% when comparing the best contextualized with the best static model (Table 5.4). GloVe embeddings outperform FastText for 27 out of 57 languages, using GloVe+PPMI (Single) boosts performance for 37 out of 57 languages, and GloVe+PPMI (All) enhances performance for 48 out of 57 languages.
XLM-R gives better performance than all static embedding configurations for only 24 out of 57 languages, but E-5 performs better than all static embeddings, showcasing some strengths of contextualized embeddings in multilingual tasks.
Cov.     ISO Contextualized Static
    E-5-B X-B     FT G GP(S) GP(A)
>95%     ro .891 .707     .405 .561 .686 .704
    sk .866 .707     .522 .67 .667 .725
    bg .869 .751     .447 .645 .711 .723
    el .872 .66     .387 .531 .712 .702
    lt .861 .704     .534 .713 .775 .797
    uk .904 .717     .542 .682 .722 .745
    lv .858 .709     .608 .737 .732 .742
    sl .848 .705     .544 .628 .715 .734
    gl .865 .723     .522 .53 .663 .699
    da .864 .724     .423 .446 .743 .717
    he .824 .701     .67 .759 .739 .784
    mk .855 .779     .598 .611 .694 .719
    ms .846 .748     .634 .694 .738 .769
[…]
>90%     et .823 .655     .583 .589 .572 .605
    be .836 .698     .674 .58 .597 .621
    az .853 .742     .667 .698 .668 .711
    eo .844 .665     .57 .504 .567 .588
    hy .809 .622     .411 .551 .609 .644
    kk .838 .69     .64 .664 .647 .69
    is .798 .651     .442 .423 .49 .534
    ka .763 .694     .591 .684 .689 .689
    ur .804 .648     .451 .42 .627 .643
    cy .704 .638     .615 .564 .608 .694
    af .865 .721     .573 .454 .56 .59
    si .809 .682     .647 .678 .613 .695
[…]
>80%     tl .85 .624     .656 .65 .707 .709
    bn .814 .595     .567 .604 .617 .681
    ga .704 .441     .585 .411 .532 .547
    mr .838 .64     .567 .627 .608 .676
    ky .783 .665     .633 .593 .571 .59
    gu .79 .613     .663 .544 .589 .631
    ml .812 .664     .641 .651 .574 .608
    pa .807 .556     .566 .474 .521 .565
    kn .803 .604     .672 .652 .581 .658
    ne .796 .699     .553 .542 .563 .605
    ha .708 .449     .421 .489 .546
    ja .802 .608     .656 .511 .523 .541
    ug .723 .606     .642 .556 .583 .622
    am .781 .559     .585 .515 .472 .555
[…]
<80%     su .765 .561     .572 .467 .446 .526
    so .642 .388     .442 .363 .403 .459
    ps .73 .542     .532 .351 .431 .493
    ht .717 .318     .531 .392 .496 .523
    yi .538 .36     .532 .341 .384 .453
    gd .54 .341     .404 .23 .397 .418
    xh .641 .324     .388 .32 .341
    yo .663 .185     .341 .199 .211 .264
    sa .762 .542     .452 .206 .261 .251
    qu .561 .245     .294 .175 .167 .153
    my .791 .564     .171 .228 .207 .163
    km .74 .631     .114 .125 .117 .109
    ku .657 .202     .09 .11 .098 .095
    lo .743 .704     .183 .185 .18
    wo .594 .238     .058 .139 .122
[…]
    All avg. .779 .591     .529 .497 .535 .566
Table 4: Macro Average F1 Scores for Topic Classification per language, sorted by GloVe vocabulary coverage. The horizontal solid lines indicate 95%, 90%, and 80% coverage by GloVe. Bold numbers indicate the maximum per line, for static and contextualized.
5.5 Additional Experiment: Graph-enhanced GloVe Improvement
To explain the improvements from injecting graph knowledge into static embeddings, we hypothesize that the size of the common vocabulary between GloVe and PPMI spaces contributes to performance gains: a larger vocabulary may lead to a better linear transformation fit, resulting in more precise projections. We investigate the relationship between common vocabulary size and embedding improvements by calculating Pearson Cohen et al. (2009) and Spearman Spearman (1961) correlations across all tasks (SID-200, SA, XNLI, SimLex).
Table 5 shows the correlation coefficients for each task and embedding configuration. For SID-200, the GloVe+PPMI (Single) embeddings have a Spearman correlation of 0.364, indicating a moderate monotonic relationship between common vocabulary count and improvement. However, the Pearson correlation of 0.096 suggests a weak linear relationship. A similar pattern can be observed in the other tasks, where Spearman correlations are generally higher than Pearson, highlighting the non-linear nature of the relationship. In contrast, tasks like SimLex show high correlations in both metrics, especially in the Single configuration, with Pearson and Spearman scores of 0.879 and 0.8, respectively.
When comparing results for Single and All configurations, the Single configurations tend to show stronger correlations. The All configurations have higher vocabulary overlaps between the embedding spaces due to contributions from various languages (C of the Appendix). This is because Single configurations focus on one language, whereas All configurations include words from multiple languages, which dilutes the strength of the relationship between vocabulary overlap and performance improvement and may suggest that the Single embedding spaces provide a better representation of graph knowledge when working in a monolingual setting.
These results suggest that while improvement scores moderately depend on vocabulary coverage, the relationship isn’t strictly linear. This implies that while a larger common vocabulary can enhance performance, other factors such as graph-based semantic knowledge may play a more significant role. Figure 1 in the Appendix visualizes the correlations for SA and SID-200 across all models.
6 Discussion: Contextual vs. Static Embeddings
While being far behind on the tested intrinsic task, the sentence embeddings extracted from E-5, a state-of-the-art multilingual sentence embedding model, consistently outperform many other configurations across all languages on the extrinsic tasks. This superior performance can be attributed to E-5’s ability to generate context-aware, sentence-level representations that capture nuanced meanings, which static embeddings, like GloVe or FastText, struggle to achieve. Unlike static word embeddings that sum individual word vectors, E-5 learns richer representations by incorporating contextual information across languages.
However, direct comparisons between E-5 and static word embeddings overlook key differences in design and use cases. E-5 is extensively trained on multilingual corpora and excels in tasks requiring complex, context-sensitive representations. In contrast, static embeddings, though simpler, are a valid alternative in low-resource or efficiency-critical scenarios, as they are effectively parameter-free during inference time. The coverage of task-specific data plays a crucial role: GloVe embeddings performe well in sentiment analysis due to broader language coverage, but poorer results in topic classification are partly linked to lower coverage in some languages. Static embeddings remain competetive across most tested extrinsic tasks and most languages, given a good vocabulary coverage.
Static embeddings enriched with external knowledge sources, such as graph-based information, provide significant advantages, especially in resource-limited applications where computational costs are critical. Computationally lightweight word vectors are invaluable in settings where models like E-5 are prohibitively expensive to deploy Strubell et al. (2019); Bommasani et al. (2021). Static embeddings also perform competitively in simpler tasks that do not heavily rely on contextual understanding Dufter et al. (2021), making them ideal for large-scale or real-time applications Gupta and Jaggi (2021). Additionally, static embeddings offer a level of transparency often lacking in complex models Vulić et al. (2020b); Bommasani et al. (2020). Their word-level semantic relationships are easy to interpret, making them useful in applications such as bias detection or model auditing.
Furthermore, Dufter et al. (2021) demonstrated that FastText outperformed BERT on a modified LAMA task Petroni et al. (2019) across ten languages while generating just 0.3% of BERT’s carbon footprint Strubell et al. (2019); Dufter et al. (2021), despite their simplicity. This highlights the overlooked value of static embeddings when evaluating resource-intensive models, rendering them useful as "green" baselines that are environmentally highly efficient.
Task P S
SID-200 (Single) 0.096 0.364
SID-200 (All) 0.284 0.115
SA (Single) 0.116 0.261
SA (All) 0.254 0.186
XNLI (Single) 0.075 0.205
XNLI (All) 0.054 0.300
SimLex (Single) 0.879 0.800
SimLex (All) 0.399 0.105
Table 5: Pearson and Spearman Correlations between Common Vocabulary Count and Improvement Scores
7 Conclusion
In this work, we developed GrEmLIn, a centralized repository of graph-enhanced GloVe embeddings for 87 mid- and low-resource languages, addressing the need for high-quality word embeddings in underrepresented languages. By merging GloVe with graph-based knowledge from ConceptNet, we enhanced the semantic richness of embeddings, leading to improved performance across tasks like semantic similarity, sentiment analysis, topic classification, and natural language inference.
Our results show that graph-enhanced GloVe outperforms the original GloVe, FastText, and even contextualized embeddings from XLM-R, offering a lightweight and environmentally efficient alternative to transformer-based models. Static embeddings have been recognized as "green" baselines, offering competitive performance at a fraction of computational cost of LLMs. This makes them ideal for low-resource settings where both computational efficiency and sustainability are key.
Limitations
While our contribution provides baseline and graph-enhanced GloVe models for many languages, several limitations exist. First, the quality and availability of training data, particularly for low-resource languages, remain key challenges. Despite leveraging large corpora like CC100 and ConceptNet, data diversity and coverage are still limited.
Second, while our method of merging GloVe embeddings with graph-based knowledge has yielded promising results, there is room for further refinement. Future work could explore more advanced fusion and projection techniques to enhance representations for low-resource languages.
Lastly, static embeddings, even with graph enhancements, cannot fully capture contextual nuances compared to transformer-based models, which may limit their performance on tasks requiring deep contextual understanding. Balancing simplicity and efficiency with improved performance remains an ongoing challenge.
Acknowledgments
We thank the anonymous reviewers for their helpful feedback. This work was supported by DisAI – Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a Horizon Europe-funded project under GA No. 101079164, and by the German Ministry of Education and Research (BMBF) as part of the project TRAILS (01IW24005).
Appendix
Appendix A Language Details
ISO code Language Size Class ConceptNet ISO code Language Size Class ConceptNet
ss Swati 86K 1 sc Sardinian 143K 1
yo Yoruba 1.1M 2 gn Guarani 1.5M 1
qu Quechua 1.5M 1 ns Northern Sotho 1.8M 1
li Limburgish 2.2M 1 ln Lingala 2.3M 1
wo Wolof 3.6M 2 zu Zulu 4.3M 2
rm Romansh 4.8M 1 ig Igbo 6.6M 1
lg Ganda 7.3M 1 as Assamese 7.6M 1
tn Tswana 8.0M 2 ht Haitian 9.1M 2
om Oromo 11M 1 su Sundanese 15M 1
bs Bosnian 18M 3 br Breton 21M 1
gd Scottish Gaelic 22M 1 xh Xhosa 25M 2
mg Malagasy 29M 1 jv Javanese 37M 1
fy Frisian 38M 0 sa Sanskrit 44M 2
my Burmese 46M 1 ug Uyghur 46M 1
yi Yiddish 51M 1 or Oriya 56M 1
ha Hausa 61M 2 la Lao 63M 2
sd Sindhi 67M 1 ta_rom Tamil Romanized 68M 3
so Somali 78M 1 te_rom Telugu Romanized 79M 1
ku Kurdish 90M 0 pu/pa Punjabi 90M 2
ps Pashto 107M 1 ga Irish 108M 2
am Amharic 133M 2 ur_rom Urdu Romanized 141M 3
km Khmer 153M 1 uz Uzbek 155M 3
bn_rom Bengali Romanized 164M 3 ky Kyrgyz 173M 3
my_zaw Burmese (Zawgyi) 178M 1 cy Welsh 179M 1
gu Gujarati 242M 1 eo Esperanto 250M 1
af Afrikaans 305M 3 sw Swahili 332M 2
mr Marathi 334M 2 kn Kannada 360M 1
ne Nepali 393M 1 mn Mongolian 397M 1
si Sinhala 452M 0 te Telugu 536M 1
la Latin 609M 3 be Belarussian 692M 3
tl Tagalog 701M 3 mk Macedonian 706M 1
gl Galician 708M 3 hy Armenian 776M 1
is Icelandic 779M 2 ml Malayalam 831M 1
bn Bengali 860M 3 ur Urdu 884M 3
kk Kazakh 889M 3 ka Georgian 1.1G 3
az Azerbaijani 1.3G 1 sq Albanian 1.3G 1
ta Tamil 1.3G 3 et Estonian 1.7G 3
lv Latvian 2.1G 3 ms Malay 2.1G 3
sl Slovenian 2.8G 3 lt Lithuanian 3.4G 3
he Hebrew 6.1G 3 sk Slovak 6.1G 3
el Greek 7.4G 3 th Thai 8.7G 3
bg Bulgarian 9.3G 3 da Danish 12G 3
uk Ukrainian 14G 3 ro Romanian 16G 3
id Indonesian 36G 3
Table 6: Details of the reproduced CC-100 corpus available on HuggingFace, including languages with their ISO codes, data set sizes, low-resource classifications, and language availability in the ConceptNet knowledge graph.
Appendix B SA Data Details
{adjustwidth}
-1cm Language ISO code Source #pos #neg #train #val #test Sundanese su Winata et al., 2023 378 383 381 76 304 Amharic am Tesfa et al., 2024 487 526 709 152 152 Swahili sw Muhammad et al., 2023a; Muhammad et al., 2023b 908 319 738 185 304 Georgian ka Stefanovitch et al., 2022 765 765 1080 120 330 Nepali ne Singh et al., 2020 680 1019 1189 255 255 Uyghur ug Li et al., 2022 2450 353 1962 311 530 Latvian lv Sproģis and Rikters, 2020 1796 1380 2408 268 500 Slovak sk Pecar et al., 2019 4393 731 3560 522 1042 Sinhala si Demotte et al., 2020 2487 2516 3502 750 751 Slovenian sl Bučar et al., 2018 1665 3337 3501 750 751 Uzbek uz Kuriyozov et al., 2019 3042 1634 3273 701 702 Bulgarian bg Martínez-García et al., 2021 6652 1271 5412 838 1673 Yoruba yo Muhammad et al., 2023a; Muhammad et al., 2023b 6344 3296 5414 1327 2899 Urdu ur Maas et al., 2011; Khan et al., 2017; Khan and Nizami, 2020 5562 5417 7356 1812 1812 Macedonian mk Jovanoski et al., 2015 3041 5184 6557 729 939 Danish da Isbister et al., 2021 5000 5000 7000 1500 1500 Marathi mr Pingle et al., 2023 5000 5000 8000 1000 1000 Bengali bn Sazzed, 2020 8500 3307 8264 1771 1772 Hebrew he Amram et al., 2018 8497 3911 8932 993 2483 Romanian ro Tache et al., 2021 7500 7500 10800 1200 3000 Telugu te Marreddy et al., 2022b; Marreddy et al., 2022a 9488 6746 11386 1634 3214 Welsh cy Espinosa-Anke et al., 2021 12500 12500 17500 3750 3750 Azerbaijani az LocalDoc, 2024 14000 14000 19600 4200 4200
Table 7: Sentiment Analysis Data Details
Appendix C Common Vocabulary Counts
ISO code GloVe and PPMI (Single) GloVe and PPMI (All)
af 9,177 85,270
am 1,105 14,217
az 7,215 80,761
be 7,623 73,750
bn 3,962 38,221
bg 92,436 368,232
ku 3,762 32,499
cy 7,774 57,522
da 38,095 450,290
el 19,710 197,647
eo 59,476 161,634
et 14,815 163,666
gd 6,415 24,430
ga 13,871 65,169
gl 29,654 215,868
gu 3,198 24,575
ht 1,557 13,304
ha 671 33,824
he 16,032 153,731
hy 14,951 60,756
is 27,007 143,567
ja 2,607 41,471
kn 2,181 24,783
ka 17,869 96,066
kk 8,292 64,494
km 2,654 34,014
ky 2,234 29,915
lo 269,010 373,012
lt 12,485 200,404
lv 17,450 183,088
ml 4,092 38,864
mr 3,211 33,552
mk 21,692 93,121
my 3,189 24,319
ne 2,650 21,479
pa 2,282 16,068
ps 847 15,904
ro 25,704 366,809
sa 3,336 12,101
si 943 27,536
sk 14,694 268,576
sl 45,153 229,429
so 533 18,088
su 1,236 26,068
sw 6,425 59,906
ta 4,596 60,906
tl 12,563 42,653
ug 764 4,798
uk 16,397 327,563
ur 4,662 44,530
uz 3,229 37,704
xh 1,650 15,709
yi 5,177 18,572
ms 34,022 152,500
yo 558 5,254
qu 2,056 11,046
wo 999 18,509
th 45,975 238,502
Table 8: Common Vocabulary between GloVe and PPMI Embedding Spaces
Appendix D Vocabulary Coverage
ISO code SA SIB XNLI MultiSimLex
G (%) F (%) G (%) F (%) G (%) F (%) G (%) F (%)
am 78.22 99.48 84.36 99.73
su 78.66 99.92 79.39 99.94
sw 88.24 100.00 91.32 99.98 83.68 99.98 73.94 100.00
si 89.18 99.99 91.63 99.97
ka 97.19 99.99 94.71 100.00
ne 77.91 99.82 84.93 99.99
ug 88.28 99.92 82.87 99.96
yo 22.37 99.18 46.50 99.73
ur 62.54 99.72 92.97 99.95 73.42 99.86
mk 82.90 99.92 95.84 99.99
mr 84.06 99.94 87.13 99.99
bn 66.55 99.75 89.58 100.00
te 85.66 99.99
uz 71.17 99.94 83.61 99.99
az 60.60 100.00 94.03 100.00
bg 84.18 99.91 98.16 100.00 96.47 99.98
sl 91.79 100.00 97.92 100.00
lv 87.04 99.41 97.43 99.97
sk 84.74 99.75 98.29 99.99
ro 90.16 99.94 98.71 100.00
he 89.72 99.74 97.57 100.00 91.79 100.00
cy 51.87 99.91 90.76 99.98 82.73 100.00
da 75.48 99.71 96.76 100.00
el 98.15 99.94 97.34 100.00
th 22.29 100.00
af 90.05 99.95
be 94.59 99.95
eo 93.83 100.00
et 94.50 100.00 94.70 99.99
gd 70.48 99.85
ga 89.97 99.93
gl 97.33 99.98
gu 87.11 99.97
ht 75.89 99.74
ha 87.20
hy 92.70 99.92
is 92.22 99.94
ja 82.97 99.98
kn 86.82 100.00
kk 93.11 100.00
km 24.06 99.92
ky 87.29 100.00
lo 19.17
lt 97.70 100.00
ml 85.13 100.00
my 31.18 99.96
pa 85.15 99.98
ps 78.72 99.91
sa 46.87 99.94
so 79.48 99.89
tl 89.07 100.00
uk 97.72 100.00
xh 62.39
yi 73.63 99.87
ms 95.71 100.00
qu 36.33 99.96
wo 54.05
Table 9: Vocabulary Coverage by GloVe FastText Embeddings for 4 Evaluation Tasks - Sentiment Analysis, Topic Classification, Natural Language Inference, and MultiSimLex
Appendix E Correlation Between Improvement Scores and Vocabulary Overlap
Figure 1: Scatter plots illustrating the relationship between vocabulary overlap and performance improvements across various language tasks using GloVe and graph-enhanced embeddings (G+P). Each plot shows the improvement in performance (G+P - GloVe) versus the common vocabulary size (log-scaled). Solid lines represent the best-fit log-linear trend.